OpenStudy (anonymous):

Solve the following equation for k. show work n= 8k-15m 3

OpenStudy (anonymous):

I got 3n+15m=k does anyone know if this is right???

OpenStudy (anonymous):

what's the "3" in the question?

OpenStudy (anonymous):

If you mean n = (8k -15m)/3 then 3n = 8k-15m 8k = 3n +15m k = 3(n+5m)/8 ) Same as you except for the 8.